OpenStudy (logan13):

A can of spaghetti has a diameter of 12 cm and a height of 16 cm. The can contains spherical meatballs, each with a radius of 3 cm, that take up 25% of the volume. How many meatballs are in the can? Use 3.14 to approximate pi and express your answer as a whole number.

OpenStudy (isaidavila):

Have choices?

OpenStudy (isaidavila):

I get 16

OpenStudy (luckycoins888):

Raduis of Can=1/2 of 12cm R=6cm Volume of Can= pi r ^{2} h = pi 6 ^{2} 16 =1809.5573cm3 Volume of Meatball=4/3 pi r ^{3} =4/3 pi 3 ^{3} =113.09734 25% of can = 1809.5573/4 =452.389325 Meatballs in area = 25%of can/volume of meatballs =452.389325/113.09734 =4 meatballs

OpenStudy (luckycoins888):

@logan13 You understand?

OpenStudy (isaidavila):

Ohh I forget 25℅

OpenStudy (luckycoins888):

LOGAN. do u understand XD

OpenStudy (anonymous):

your welcome!

OpenStudy (luckycoins888):

XD thx lmao

OpenStudy (luckycoins888):

logan are u ded or something